[figure: 9.7 cm]

APPLICATION OF MAGNETICALLY MODULATED MICROWAVE ABSORPTION TO STUDY GMR EFFECT IN

Ni-Fe/Cu MULTILAYER SYSTEM

J. Stankowski, F. Stobiecki, M. Górska Institute of Molecular Physics, Polish Academy of Sciences Smoluchowskiego 17, PL-60179 Pozna, Poland

Magnetically Modulated Microwave Absorption (MMMA), Magnetic Hysteresis M(H), Giant Magnetoresistance (GMR) and Ferromagnetic Resonance (FMR) were examined in antiferromagnetically coupled Py/Cu (Py=Ni_{83}Fe_{17}-permalloy) multilayers (ML’s). The correlation between MMMA technique and the standard GMR measurements as well as Magnetisation Reversal process will be showed. This work presents microwave studies of GMR, MR and FMR effects for different orientation of magnetic field direction relative to the sample surface. In Py/Cu ML’s, at zero magnetic field, two component MMMA signal: broad (35 Oe) and narrow (2 Oe), has been observed. The broad line is related to the GMR effect. The extreme points of the MMMA broad line correspond to the maximum slope of the GMR(H) function. The narrow MMMA line is due to the magnetisation reversal MR, extremes of that are related to coercive field values H_{c}. MMMA represents a new contactless method to study GMR effect in CIP configuration. This technique could be also used to determine very small coercive field values H_{c} in magnetic materials.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : J. Stankowski

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences Smoluchowskiego 17, PL-60179 Pozna, Poland

Email address : Jan.Stankowski@ifmpan.poznan.pl